# The Ginsburg–Sands theorem and computability theory

### v1

@@ 5. Hausdorff spaces
If it exists, fix the strongest requirement with index smaller than $s$ that requires attention at stage $s$. If this is $\mathcal{D}_{x}$ for some $x$, choose the least $y$ so that $12\langle x,y\rangle+$ 1 is larger than all indices claimed by any requirement at any point in the construction, including claims that may since have been cancelled. Since at most one requirement claims at most finitely many numbers at any stage, $y$ exists. Now say that $\mathcal{D}_{x}$ claims $12\langle x,y\rangle+$ 1. If the requirement is instead $\mathcal{R}_{e,u}$ for some $e$ and $u$, fix the least witnessing $x,\langle n_{0},\ldots,n_{r}\rangle<s$. Cancel the claims of any lower priority requirements, and say that $\mathcal{R}_{e,u}$ claims $n_{0},\ldots,n_{r}$. In either case, say that the requirement in question acts. Finally, for each $n$ that is claimed by some $\mathcal{R}$-requirement at this point, enumerate $s$ into $V_{n}$. This completes the construction. Since the only time a number $s$ is enumerated into some $V_{n}$ during the construction is at stage $s$, the sequence $\langle V_{n}:n\in\omega\rangle$ is computable. It follows that the CSC space $\langle\omega,\mathcal{U},k\rangle$ generated from $\langle V_{n}:n\in\omega\rangle$ is computable as well. Next, note that if $n$ is even and $n\neq\overline{n}$, then $n$ and $\overline{n}$ can never be claimed at the same time (whether by the same requirement, or by different ones). Indeed, even indices $n$ can only be claimed by $\mathcal{R}$-requirements. When an even $n$ is claimed by a new $\mathcal{R}$-requirement at some stage $s$, then $\overline{n}$ must not already be claimed by any higher priority requirement (necessarily another $\mathcal{R}$-requirement). And if $\overline{n}$ was already claimed by a lower priority requirement, then this claim is first cancelled. It follows that, at the end of each stage $s$, the number $s$ is enumerated into at most one of $V_{n}$ or $V_{\overline{n}}$ for every even $n$ with $n\neq\overline{n}$. In particular, if $x\neq y$ then $V_{2\langle x,y\rangle}$ and $V_{2\langle y,x\rangle}$ are disjoint. By our remarks above, $\langle\omega,\mathcal{U},k\rangle$ is effectively Hausdorff. An induction on the priority ordering shows that each requirement acts only finitely often. This is because once a requirement acts at a stage after all stronger $\mathcal{R}$-requirements have stopped acting, it claims some indices, and these claims cannot later be cancelled. Hence, the requiurement can never require attention and act again. It thus remains only to show that all requirements are satisfied. Fix a requirement, and let $s_{0}$ be least such that every stronger requirement has stopped acting prior to stage $s_{0}$. Thus, either $s_{0}=0$ or some stronger $\mathcal{R}$-requirement acted at stage $s_{0}-1$ and thereby cancelled any claim the requirement in question may have had. Either way, then, this requirement has no claims at the start of stage $s_{0}$. Hence, if the requirement in question is $\mathcal{D}_{x}$ for some $x$, then it requires attention at the least stage $s\geq s_{0}$ with $s>x$, at which point it claims some $12\langle x,y\rangle+$ 1 that has never been claimed by any requirement before. By choice of $s_{0}$, this claim is never cancelled at any stage after $s$, and so in particular, $12\langle x,y\rangle+$ 1 is never claimed by any $\mathcal{R}$-requirement. It follows that no numbers are enumerated into $V_{2\langle x,y\rangle+1}$ during the construction, so $V_{2\langle x,y\rangle+1}$ is just the singleton $\{x\}$ and $\mathcal{D}_{x}$ is satisfied. Now suppose the requirement in question is $\mathcal{R}_{e,u}$ for some $e,u$. Seeking a contradiction, suppose $\Phi_{e}$ is the characteristic function of an infinite set $Y$, $\Phi_{u}$ is total on $Y$, and for all $x\in Y$ we have $x\in\bigcap_{i\leq r}V_{n_{i}}$ where $\Phi_{u}(x)\operatorname{\downarrow}=\langle n_{0},\ldots,n_{r}\rangle$. Let $F_{0}$ be the set of all indices $n$ claimed by stronger $\mathcal{D}$-requirement prior to stage $s_{0}$, and $F_{1}$ the set of all indices $n$ claimed by stronger $\mathcal{R}$-requirements. By choice of $s_{0}$, none of these requirements acts at any stage $s\geq s_{0}$, so their claims in $F_{0}\cup F_{1}$ are permanent and none of them claims any other numbers. By our earlier observations, if $n\in F_{0}$ then $V_{n}$ is a singleton, and if $n\in F_{1}$ is even and $n\neq\overline{n}$ then $V_{n}$ and $V_{\overline{n}}$ are disjoint. Moreover, by construction, if $n\in F_{1}$ then $V_{n}$ is cofinite, so if $n$ is even and $n\neq\overline{n}$ then $V_{\overline{n}}$ is finite. Let $m$ be the maximum of all $V_{n}$ for $n\in F_{0}$ and all $V_{\overline{n}}$ for even $n\in F_{1}$ with $n\neq\overline{n}$. Since $Y$ is infinite, we can choose $x>\max\{m,s_{0}\}$ in $Y$. Then if $\Phi_{u}(x)\operatorname{\downarrow}=\langle n_{0},\ldots,n_{r}\rangle$ we have $x\in V_{n_{i}}$ for all $i\leq r$, and hence it cannot be that $n_{i}\in F_{0}$ or $\overline{n_{i}}\in F_{1}$ if $n_{i}$ is even and $n_{i}\neq\overline{n_{i}}$. We conclude that, if we take the least $s\geq s_{0}$ for which there exists $x,\langle n_{0},\ldots,n_{r}\rangle<s$ with these properties, and such that $\Phi_{e}(x)[s]\operatorname{\downarrow}=1$ and $\Phi_{u}(x)[s]\operatorname{\downarrow}=\langle n_{0},\ldots,n_{r}\rangle$, then $\mathcal{R}_{e,u}$ will require attention at stage $s$. It will then claim $n_{0},\ldots,n_{r}$, and by choice of $s_{0}$, these claims will be permanent. Hence, $V_{n_{i}}$ will be cofinite for all $i\leq r$, proving that $\bigcap_{i\leq r}V_{n_{i}}$ cannot be $\{x\}$ after all. The proof is complete. ∎
@@ 6.1. An alternative proof of $\mathsf{GS}$ for $T_{1}$ CSC spaces
To prove (1), fix any $x\in U\in\mathcal{U}$. Since $X$ is $T_{1}$, for each $y\neq x$ there is a $V\in\mathcal{U}$ containing $y$ but not $x$. Now by Proposition 3.7, if $U$ were finite we would have $\{x\}\in\mathcal{U}$, whence $x$ and $y$ would be a $T_{2}$ pair as witnessed by $\{x\}$ and $V$, which cannot be. To prove (2), fix a finite set $F=\{n_{0}<\cdots<n_{r}\}$ and suppose $U_{n_{i}}\neq\emptyset$ for all $i\leq r$. We use $\Sigma^{0}_{1}$ induction on $j\leq r$ to show that $\bigcap_{i\leq j}U_{n_{i}}\neq\emptyset$. This is clear for $j=0$. Suppose it is true for some $j<r$, fix $x\in\bigcap_{i\leq j}U_{n_{i}}$, and let $n=\overline{k}(x,\{n_{i}:i\leq j\})$, where $\overline{k}$ is as given by Lemma 3.2. Then $x\in U_{n}\subseteq\bigcap_{i\leq j}U_{n_{i}}$. Now if $U_{n}\cap U_{n_{j+1}}$ were empty, then every element of $U_{n}$ would form a $T_{2}$ pair with every element of $U_{n_{j+1}}$, which is impossible. Hence, $U_{n}\cap U_{n_{j+1}}\subseteq\bigcap_{i\leq{j+1}}U_{n_{i}}$ is nonempty, as desired. We conclude that $\bigcap_{n\in F}U_{n}\neq\emptyset$. Fix $x$ in this intersection. Then $x\in U_{\overline{k}(x,F)}\subseteq\bigcap_{n\in F}U_{n}\neq\emptyset$. By (1), $U_{\overline{k}(x,F)}$ is infinite, so $\bigcap_{n\in F}U_{n}$ is too. To prove (3), we proceed as follows. Write $\mathcal{U}=\langle U_{n}:n\in\mathbb{N}\rangle$. $\mathsf{RCA}_{0}$ proves that there exists a function $f:\mathbb{N}\to\mathbb{N}$ such that $\operatorname{ran}(f)=\{n:U_{n}\neq\emptyset\}$. By (2), $\bigcap_{i\leq j}U_{f(i)}$ is infinite for every $j$. Let $x_{0}$ be any element of $U_{f(0)}$, and having defined $x_{i}$ for all $i<j$, let $x_{j}$ be the least element of $\bigcap_{i\leq j}U_{f(i)}$ larger than all the $x_{i}$. Let $Y=\{x_{i}:i\in\omega\}$. Then for each nonempty $U_{n}$, say with $f(i)=n$, we have that $x_{j}\in U_{f(i)}=U_{n}$ for all $j\geq i$. Thus, $Y$ is an infinite computable subspace of $X$ with the cofinite topology, as was to be shown. ∎
@@ 6.2. $\mathsf{GST}_{1}$ and $\mathsf{RT}^{2}_{2}$
We argue in $\mathsf{RCA}_{0}$. Fix an infinite linear order $\langle L,\leq_{L}\rangle$. We define a $T_{1}$ CSC space with the property that any infinite discrete subspace is an ascending sequence under $\leq_{L}$, and any infinite subspace with the cofinite topology is a descending sequence. This yields $\mathsf{ADS}$. For each $x\in L$, let
@@
Let $\langle L,\mathcal{U},k\rangle$ be the CSC space on $L$ generated by $\langle V_{n}:n\in\mathbb{N}\rangle$. Note that $X$ is $T_{1}$: if $x\neq y$, then $V_{\langle x,y\rangle}$ contains $x$ but not $y$, and $V_{\langle y,x\rangle}$ contains $y$ but not $x$. Now, consider any infinite subspace $Y$ of $X$ with either the discrete or cofinite topology. We consider the two cases separately. If $Y$ has the discrete topology, then for every $x\in Y$ there is a $U\in\mathcal{U}$ such that $U\cap Y=\{x\}$. Then there is a finite set $F$ such that $U=\bigcap_{y\in F}V_{\langle x,y\rangle}$. (For if $x\in V_{\langle z,y\rangle}$ for some $z$ then necessarily $x\leq_{L}z$ and so $V_{\langle x,y\rangle}\subseteq V_{\langle z,y\rangle}$.) It follows that |$|\{w\in Y:w\leq_{L}x\}|\leq|F|$, so $x$ has only finitely many $\leq_{L}$-predecessors in $Y$. Hence, $Y$ is an ascending sequence under $\leq_{L}$. If $Y$ has the cofinite topology, we have instead that for every $x\in Y$ and every $U\in\mathcal{U}$, if $x\in U$ then $U\cap Y=^{*}Y$. In particular, $V_{\langle x,x\rangle}\cap Y=^{*}Y$, meaning that almost every $z\in Y$ satisfies $z\leq_{L}x$. In other words, $x$ has only finitely many $\leq_{L}$-successors in $Y$. We conclude that $Y$ is a descending sequence under $\leq_{L}$, as was to be shown. ∎
@@
As discussed in Section 1, $\mathsf{ADS}$ implies $\mathsf{COH}$ by [20, Proposition 2.10], so by the preceding theorem $\mathsf{GST}_{1}$ implies $\mathsf{COH}$. Since $\mathsf{GST}_{1}$ implies $\mathsf{SGST}_{1}$, the result now follows by Proposition 6.4. ∎
@@
    $(\exists Y)[Y\text{ is infinite }\wedge((\forall x)[x\in Y\leftrightarrow\varphi(x)]\vee(\forall x)[x\in Y\leftrightarrow\neg\varphi(x)])],$
@@
  • $\Delta^{0}_{2}\text{-}\mathsf{Subset}$ is the scheme consisting of all formulas of the form
@@
    $(\exists Y)[Y\text{ is infinite }\wedge((\forall x)[x\in Y\leftrightarrow\varphi(x)]\vee(\forall x)[x\in Y\leftrightarrow\neg\varphi(x)])],$
@@
and observe the following fact about all $x$ and $y$. If $(\forall z)\theta(x,y,z)$ holds, then $V_{\langle x,y,s\rangle}=\{x\}$ for all $s$. Otherwise, there is a least $z$ such that $\neg\theta(x,y,z)$ holds since $\theta$ is $\Delta^{0}_{0}$, and $V_{\langle x,y,s\rangle}=\{x\}\cup\{w>\max\{x,s\}:w\geq z\}$ for all $s$. In particular, for all $x$ and $y$ either $V_{\langle x,y,s\rangle}=\{x\}$ for all $s$, or $V_{\langle x,y,s\rangle}$ is a cofinite subset of $\mathbb{N}$ for all $s$. Moreover, there is a $y$ such that $V_{\langle x,y,s\rangle}=\{x\}$ for some (equivalently, all) $s$ if and only if $\varphi(x)$ holds. Let $\langle\mathbb{N},\mathcal{U},k\rangle$ be the CSC space generated by $\langle V_{n}:n\in\mathbb{N}\rangle$. Fix any $x\in\mathbb{N}$ such that $\neg\varphi(x)$ holds, and any $U\in\mathcal{U}$ such that $x\in U$. We claim that $U$ is cofinite. Write $U=\bigcap_{n\in F}V_{n}$ for some nonempty finite set $F$, and for each $n\in F$ write $n=\langle x_{n},y_{n},s_{n}\rangle$. If $x_{n}=x$, then $V_{\langle x_{n},y_{n},s_{n}\rangle}$ is cofinite by the argument above because $\neg\varphi(x)$ holds. If $x_{n}\neq x$ then $V_{\langle x_{n},y_{n},s_{n}\rangle}\neq\{x_{n}\}$ since it contains $x$, hence it is also cofinite. Thus, each $V_{n}$ for $n\in F$ is cofinite. For each $n\in F$, fix the least $z_{n}>x_{n}$ in $V_{n}$; the set $D=\{z_{n}:n\in F\}$ exists by $\Delta^{0}_{1}$ comprehension and is nonempty. Furthermore, as observed above, $V_{n}=\{x_{n}\}\cup\{w:w\geq z_{n}\}$. Hence, every $w\geq\max D$ belongs to $\bigcap_{n\in F}V_{n}=U$, which proves the claim. We conclude that $X$ is stable: if $\varphi(x)$ holds then $V_{\langle x,y,0\rangle}=\{x\}$ for any $y$ such that $(\exists y)(\forall z)\theta(x,y,z)$ holds; if $\neg\varphi(x)$ then every $U\in\mathcal{U}$ containing $x$ is cofinite. Moreover, $X$ is $T_{1}$: given $x\neq y$, we have that $V_{x,0,y}$ contains $x$ but not $y$, while $V_{y,0,x}$ contains $y$ but not $x$. Let $Y$ be any infinite subspace of $X$ with either the discrete or cofinite topology. By construction, in the former case we have that $\varphi(x)$ holds for all $x\in Y$; in the latter case, we have that $\neg\varphi(x)$ holds for all $x\in Y$. Hence, $\Sigma^{0}_{2}$-$\mathsf{Subset}$ holds. Conversely, assume $\Sigma^{0}_{2}$-$\mathsf{Subset}$ and let $\langle X,\mathcal{U},k\rangle$ be any infinite stable $T_{1}$ CSC space. Write $\mathcal{U}=\langle U_{n}:n\in\mathbb{N}\rangle$, and let $\varphi(x)$ be the $\Sigma^{0}_{2}$ formula $(\exists n)[U_{n}=\{x\}]$. Notice that by stability of $X$, $\neg\varphi(x)$ holds if and only if every $U\in\mathcal{U}$ containing $x$ is cofinite. Let $Y$ be any infinite set such that either $\varphi(x)$ holds for all $x\in Y$, or $\neg\varphi(x)$ holds for all $x\in Y$. In the former case, we have that $Y$ forms a discrete subspace of $X$; in the latter, it is a subspace with the cofinite cofinite topology. So, $\mathsf{SGST}_{1}$ holds. ∎

### v2

@@ 5. Hausdorff spaces
Now suppose the requirement in question is $\mathcal{R}_{e,u}$ for some $e,u$. Seeking a contradiction, suppose $\Phi_{e}$ is the characteristic function of an infinite set $Y$, $\Phi_{u}$ is total on $Y$, and for all $x\in Y$ we have $x\in\bigcap_{i\leq r}V_{n_{i}}$ where $\Phi_{u}(x)\operatorname{\downarrow}=\langle n_{0},\ldots,n_{r}\rangle$. Let $F_{0}$ be the set of all indices $n$ claimed by stronger $\mathcal{D}$-requirement prior to stage $s_{0}$, and $F_{1}$ the set of all indices $n$ claimed by stronger $\mathcal{R}$-requirements. By choice of $s_{0}$, none of these requirements acts at any stage $s\geq s_{0}$, so their claims in $F_{0}\cup F_{1}$ are permanent and none of them claims any other numbers. By our earlier observations, if $n\in F_{0}$ then $V_{n}$ is a singleton, and if $n\in F_{1}$ is even and $n\neq\overline{n}$ then $V_{n}$ and $V_{\overline{n}}$ are disjoint. Moreover, by construction, if $n\in F_{1}$ then $V_{n}$ is cofinite, so if $n$ is even and $n\neq\overline{n}$ then $V_{\overline{n}}$ is finite. Let $m$ be the maximum of all $V_{n}$ for $n\in F_{0}$ and all $V_{\overline{n}}$ for even $n\in F_{1}$ with $n\neq\overline{n}$. Since $Y$ is infinite, we can choose $x>\max\{m,s_{0}\}$ in $Y$. Then if $\Phi_{u}(x)\operatorname{\downarrow}=\langle n_{0},\ldots,n_{r}\rangle$ we have $x\in V_{n_{i}}$ for all $i\leq r$, and hence it cannot be that $n_{i}\in F_{0}$ or $\overline{n_{i}}\in F_{1}$ if $n_{i}$ is even and $n_{i}\neq\overline{n_{i}}$. We conclude that, if we take the least $s\geq s_{0}$ for which there exists $x,\langle n_{0},\ldots,n_{r}\rangle<s$ with these properties, and such that $\Phi_{e}(x)[s]\operatorname{\downarrow}=1$ and $\Phi_{u}(x)[s]\operatorname{\downarrow}=\langle n_{0},\ldots,n_{r}\rangle$, then $\mathcal{R}_{e,u}$ will require attention at stage $s$. It will then claim $n_{0},\ldots,n_{r}$, and by choice of $s_{0}$, these claims will be permanent. Hence, $V_{n_{i}}$ will be cofinite for all $i\leq r$, proving that $\bigcap_{i\leq r}V_{n_{i}}$ cannot be $\{x\}$ after all. The proof is complete. ∎
@@ 6.1. An alternative proof of $\mathsf{GS}$ for $T_{1}$ CSC spaces
To prove (1), fix any $x\in U\in\mathcal{U}$. Since $X$ is $T_{1}$, for each $y\neq x$ there is a $V\in\mathcal{U}$ containing $y$ but not $x$. Now by Proposition 3.7, if $U$ were finite we would have $\{x\}\in\mathcal{U}$, whence $x$ and $y$ would be a $T_{2}$ pair as witnessed by $\{x\}$ and $V$, which cannot be.
To prove (2), fix a finite set $F=\{n_{0}<\cdots<n_{r}\}$ and suppose $U_{n_{i}}\neq\emptyset$ for all $i\leq r$. We use $\Sigma^{0}_{1}$ induction on $j\leq r$ to show that $\bigcap_{i\leq j}U_{n_{i}}\neq\emptyset$. This is clear for $j=0$. Suppose it is true for some $j<r$, fix $x\in\bigcap_{i\leq j}U_{n_{i}}$, and let $n=\overline{k}(x,\{n_{i}:i\leq j\})$, where $\overline{k}$ is as given by Lemma 3.2. Then $x\in U_{n}\subseteq\bigcap_{i\leq j}U_{n_{i}}$. Now if $U_{n}\cap U_{n_{j+1}}$ were empty, then every element of $U_{n}$ would form a $T_{2}$ pair with every element of $U_{n_{j+1}}$, which is impossible. Hence, $U_{n}\cap U_{n_{j+1}}\subseteq\bigcap_{i\leq{j+1}}U_{n_{i}}$ is nonempty, as desired. We conclude that $\bigcap_{n\in F}U_{n}\neq\emptyset$. Fix $x$ in this intersection. Then $x\in U_{\overline{k}(x,F)}\subseteq\bigcap_{n\in F}U_{n}\neq\emptyset$. By (1), $U_{\overline{k}(x,F)}$ is infinite, so $\bigcap_{n\in F}U_{n}$ is too.
To prove (3), we proceed as follows. Write $\mathcal{U}=\langle U_{n}:n\in\mathbb{N}\rangle$. $\mathsf{RCA}_{0}$ proves that there exists a function $f:\mathbb{N}\to\mathbb{N}$ such that $\operatorname{ran}(f)=\{n:U_{n}\neq\emptyset\}$. By (2), $\bigcap_{i\leq j}U_{f(i)}$ is infinite for every $j$. Let $x_{0}$ be any element of $U_{f(0)}$, and having defined $x_{i}$ for all $i<j$, let $x_{j}$ be the least element of $\bigcap_{i\leq j}U_{f(i)}$ larger than all the $x_{i}$. Let $Y=\{x_{i}:i\in\omega\}$. Then for each nonempty $U_{n}$, say with $f(i)=n$, we have that $x_{j}\in U_{f(i)}=U_{n}$ for all $j\geq i$. Thus, $Y$ is an infinite computable subspace of $X$ with the cofinite topology, as was to be shown. ∎
@@ 6.2. $\mathsf{GST}_{1}$ and $\mathsf{RT}^{2}_{2}$
We argue in $\mathsf{RCA}_{0}$. Fix an infinite linear order $\langle L,\leq_{L}\rangle$. We define a $T_{1}$ CSC space with the property that any infinite discrete subspace is an ascending sequence under $\leq_{L}$, and any infinite subspace with the cofinite topology is a descending sequence. This yields $\mathsf{ADS}$.
For each $x\in L$, let
@@
Let $\langle L,\mathcal{U},k\rangle$ be the CSC space on $L$ generated by $\langle V_{n}:n\in\mathbb{N}\rangle$. Note that $X$ is $T_{1}$: if $x\neq y$, then $V_{\langle x,y\rangle}$ contains $x$ but not $y$, and $V_{\langle y,x\rangle}$ contains $y$ but not $x$.
Now, consider any infinite subspace $Y$ of $X$ with either the discrete or cofinite topology. We consider the two cases separately.
If $Y$ has the discrete topology, then for every $x\in Y$ there is a $U\in\mathcal{U}$ such that $U\cap Y=\{x\}$. Then there is a finite set $F$ such that $U=\bigcap_{y\in F}V_{\langle x,y\rangle}$. (For if $x\in V_{\langle z,y\rangle}$ for some $z$ then necessarily $x\leq_{L}z$ and so $V_{\langle x,y\rangle}\subseteq V_{\langle z,y\rangle}$.) It follows that |$|\{w\in Y:w\leq_{L}x\}|\leq|F|$, so $x$ has only finitely many $\leq_{L}$-predecessors in $Y$. Hence, $Y$ is an ascending sequence under $\leq_{L}$.
If $Y$ has the cofinite topology, we have instead that for every $x\in Y$ and every $U\in\mathcal{U}$, if $x\in U$ then $U\cap Y=^{*}Y$. In particular, $V_{\langle x,x\rangle}\cap Y=^{*}Y$, meaning that almost every $z\in Y$ satisfies $z\leq_{L}x$. In other words, $x$ has only finitely many $\leq_{L}$-successors in $Y$. We conclude that $Y$ is a descending sequence under $\leq_{L}$, as was to be shown. ∎
@@
As discussed in Section 1, $\mathsf{ADS}$ implies $\mathsf{COH}$ by [21, Proposition 2.10], so by the preceding theorem $\mathsf{GST}_{1}$ implies $\mathsf{COH}$. Since $\mathsf{GST}_{1}$ implies $\mathsf{SGST}_{1}$, the result now follows by Proposition 6.4. ∎
@@
    $(\exists Y)[Y\text{ is infinite }\wedge((\forall x)[x\in Y\to\varphi(x)]\vee(\forall x)[x\in Y\to\neg\varphi(x)])],$
@@
  • $\Delta^{0}_{n}\text{-}\mathsf{Subset}$ is the scheme consisting of all formulas of the form
@@
    $(\exists Y)[Y\text{ is infinite }\wedge((\forall x)[x\in Y\to\varphi(x)]\vee(\forall x)[x\in Y\to\neg\varphi(x)])],$
@@
and observe the following fact about all $x$ and $y$. If $(\forall z)\theta(x,y,z)$ holds, then $V_{\langle x,y,s\rangle}=\{x\}$ for all $s$. Otherwise, there is a least $z$ such that $\neg\theta(x,y,z)$ holds since $\theta$ is $\Delta^{0}_{0}$, and $V_{\langle x,y,s\rangle}=\{x\}\cup\{w>\max\{x,s\}:w\geq z\}$ for all $s$. In particular, for all $x$ and $y$ either $V_{\langle x,y,s\rangle}=\{x\}$ for all $s$, or $V_{\langle x,y,s\rangle}$ is a cofinite subset of $\mathbb{N}$ for all $s$. Moreover, there is a $y$ such that $V_{\langle x,y,s\rangle}=\{x\}$ for some (equivalently, all) $s$ if and only if $\varphi(x)$ holds.
Let $\langle\mathbb{N},\mathcal{U},k\rangle$ be the CSC space generated by $\langle V_{n}:n\in\mathbb{N}\rangle$. Fix any $x\in\mathbb{N}$ such that $\neg\varphi(x)$ holds, and any $U\in\mathcal{U}$ such that $x\in U$. We claim that $U$ is cofinite. Write $U=\bigcap_{n\in F}V_{n}$ for some nonempty finite set $F$, and for each $n\in F$ write $n=\langle x_{n},y_{n},s_{n}\rangle$. If $x_{n}=x$, then $V_{\langle x_{n},y_{n},s_{n}\rangle}$ is cofinite by the argument above because $\neg\varphi(x)$ holds. If $x_{n}\neq x$ then $V_{\langle x_{n},y_{n},s_{n}\rangle}\neq\{x_{n}\}$ since it contains $x$, hence it is also cofinite. Thus, each $V_{n}$ for $n\in F$ is cofinite. For each $n\in F$, fix the least $z_{n}>x_{n}$ in $V_{n}$; the set $D=\{z_{n}:n\in F\}$ exists by $\Delta^{0}_{1}$ comprehension and is nonempty. Furthermore, as observed above, $V_{n}=\{x_{n}\}\cup\{w:w\geq z_{n}\}$. Hence, every $w\geq\max D$ belongs to $\bigcap_{n\in F}V_{n}=U$, which proves the claim.
We conclude that $X$ is stable: if $\varphi(x)$ holds then $V_{\langle x,y,0\rangle}=\{x\}$ for any $y$ such that $(\exists y)(\forall z)\theta(x,y,z)$ holds; if $\neg\varphi(x)$ then every $U\in\mathcal{U}$ containing $x$ is cofinite. Moreover, $X$ is $T_{1}$: given $x\neq y$, we have that $V_{x,0,y}$ contains $x$ but not $y$, while $V_{y,0,x}$ contains $y$ but not $x$. Let $Y$ be any infinite subspace of $X$ with either the discrete or cofinite topology. By construction, in the former case we have that $\varphi(x)$ holds for all $x\in Y$; in the latter case, we have that $\neg\varphi(x)$ holds for all $x\in Y$. Hence, $\Sigma^{0}_{2}$-$\mathsf{Subset}$ holds.
Conversely, assume $\Sigma^{0}_{2}$-$\mathsf{Subset}$ and let $\langle X,\mathcal{U},k\rangle$ be any infinite stable $T_{1}$ CSC space. Write $\mathcal{U}=\langle U_{n}:n\in\mathbb{N}\rangle$, and let $\varphi(x)$ be the $\Sigma^{0}_{2}$ formula $(\exists n)[U_{n}=\{x\}]$. Notice that by stability of $X$, $\neg\varphi(x)$ holds if and only if every $U\in\mathcal{U}$ containing $x$ is cofinite. Let $Y$ be any infinite set such that either $\varphi(x)$ holds for all $x\in Y$, or $\neg\varphi(x)$ holds for all $x\in Y$. In the former case, we have that $Y$ forms a discrete subspace of $X$; in the latter, it is a subspace with the cofinite cofinite topology. So, $\mathsf{SGST}_{1}$ holds. ∎